OpenStudy (anonymous):

tell whether the terms are like terms 3a and -5a

OpenStudy (anonymous):

i have another question how can i rewrite this expression as a sum -8x2 + 3xy - 9x - 3

OpenStudy (anonymous):

the 2 is an exponent .

OpenStudy (cgreenwade2000):

It's already a sum.

OpenStudy (phi):

write -8x^2 to show an exponent. notice that all the terms are different: different letters or different powers

OpenStudy (anonymous):

i dont get it what is the sum ? and so 3a and -5a is a like terms ?

OpenStudy (phi):

same letter to the same power is a "like term" 3a and -5a are "like terms" a^1 or simply a so you can add: 3 a's take away 5 a's = -2a

OpenStudy (anonymous):

also 2ab-5ab^2 -9a^2b

OpenStudy (phi):

as for 2ab-5ab^2 -9a^2b These are all different terms, so that is as simple as it gets ab would match ab or ba ab^2 would match ab^2 or (b^2)a (order does not matter) same for a^2 b

OpenStudy (anonymous):

rewrite each expression as a sum "-8x^2+3xy-9x-3" "2ab-5ab^2-9a^2b" how do i rewrite these as a sum ??

OpenStudy (phi):

they already are sums. if you want to get fussy, you might write them alphabetically, highest order term first the first expression is already in that order . the 2nd might be written as -9a^2b -5a b^2 +2 a b but that is all you can do, and is not really necessary.

OpenStudy (anonymous):

how can you express 499 to find the product 499 x 5 using mental math ? Explain.

OpenStudy (anonymous):

thank you for all this help i suck in math :/

OpenStudy (cgreenwade2000):

What easy number to multiply is very close to 499?

OpenStudy (phi):

I would say 499 is almost 500 so 500*5= 2500 then subtract the 5*1 to get 2495

OpenStudy (phi):

in other words (500-1)*5 is the same problem use the distributive law to get : 500*5 -5= 5*5 * 100 - 5